OpenStudy (teller):

Would anyone mind helping me to see if I did this correctly? At the city museum, child admission is $5.90 and adult admission is $9.00.On Sunday, three times as many adult tickets as child tickets were sold, for a total sales of $921.20.How many child tickets were sold that day? C+A=14.9 5.90C+9.00A=921.20 5.90(14.9-A)+9.00A=921.20 5.90-14.9=87.91 87.91-5.90+9.00A=921.20 -3.1A=833.29 CANCEL OUT THE -3.1 AND DIVIDE 833.29 BY -3.1A =268.80 SO WOULD IT BE A=268.80?

Vocaloid (vocaloid):

uh, not quite, remember that A = the number of adult tickets, so it should be a whole number (no decimals!) how about we start over from the very beginning, ok?

Vocaloid (vocaloid):

ok, we let C = number of child's tickets and A = number of adults tickets, with me so far?

OpenStudy (teller):

there should be a decimal since it is money right?

Vocaloid (vocaloid):

think about what A and C represent...

Vocaloid (vocaloid):

A = number of adult tickets C = number of child tickets there's no such thing as a half a ticket, correct? the number of tickets must be a whole number

OpenStudy (teller):

Um okay

Vocaloid (vocaloid):

well, the problem tells us that "three times as many adult tickets as child tickets were sold" which tells us that: A = 3*C, agreed?

OpenStudy (teller):

Yes

Vocaloid (vocaloid):

and we know that (price of adult ticket)(number of adult tickets) + (price of child's ticket)(number of child's tickets) = total sales

OpenStudy (teller):

alright

Vocaloid (vocaloid):

giving us: 5.90C + 9.00A = 921.20, correct?

OpenStudy (teller):

Yes

OpenStudy (teller):

so i had that part right

Vocaloid (vocaloid):

now let's work w/ our two equations: A = 3C 5.90C + 9.00A = 921.20

OpenStudy (teller):

wait let me gues

OpenStudy (teller):

5.90(3-a)+9.00a=921.20?

Vocaloid (vocaloid):

almost but not quite...

OpenStudy (teller):

oh lol

Vocaloid (vocaloid):

A = 3C, so we can replace "A" with "3C" whenever A shows up

Vocaloid (vocaloid):

giving us: 5.90C + 9.00A = 921.20 5.90C + 9.00(3C) = 921.20

OpenStudy (teller):

alright

Vocaloid (vocaloid):

now let's work on the left side of the equation... 5.90C + 9.00(3C) = 5.90C + 27C = 32.9C, agreed?

OpenStudy (teller):

yes

Vocaloid (vocaloid):

so that leaves us with 32.9C = 921.20 do you know how to find C from here?

OpenStudy (teller):

would I divide ?

Vocaloid (vocaloid):

yup, divide both sides by 32.9 to find C

OpenStudy (teller):

I got an odd number...

Vocaloid (vocaloid):

check your math again, use a calculator if you aren't sure

Vocaloid (vocaloid):

921.20/32.9 = ?

OpenStudy (teller):

0.0357142857142857

Vocaloid (vocaloid):

other way around, big number divided by small number 921.20/32.9 = ?

OpenStudy (teller):

28

OpenStudy (teller):

ooooo thank you

Vocaloid (vocaloid):

right, so C = 28, and since C is the number of child's tickets, 28 is our answer

OpenStudy (teller):

thank you!

OpenStudy (teller):

@Vocaloid mind helping me with another one it's my last question

Vocaloid (vocaloid):

sure, I'll give it a shot

Vocaloid (vocaloid):

post a new question and tag me